OpenStudy (anonymous):

x+2 / x^2 -9 + x-3/x^2-2x-3 = 2x-3/ x^2+4x+3

OpenStudy (anonymous):

what about it?

OpenStudy (anonymous):

algebra , i dont know how to solve it.

OpenStudy (anonymous):

kathyworld, it is not clear how the equation is formed. Can you please put in some brackets or parentheses to make the equation better?

OpenStudy (anonymous):

[ x+2 / x^2 -9] + [x-3/x^2-2x-3] = 2x-3/ x^2+4x+3

OpenStudy (anonymous):

^ = power

OpenStudy (anonymous):

I assume it looks like this: [(x+2 )/( x^2 -9)] + [(x-3)/(x^2-2x-3)] = (2x-3)/ (x^2+4x+3)

OpenStudy (anonymous):

yes perfect :)

OpenStudy (anonymous):

ok, thanks

OpenStudy (anonymous):

i just dont understand how to solve it

OpenStudy (anonymous):

[(x+2) / (x-3)(x+3)] + [(x-3) / (x-3)(x+1)] = (2x-3)/(x+3)(x+1)

OpenStudy (anonymous):

is that the answer ?

OpenStudy (anonymous):

[(x+2 )/( x^2 -9)] + [(x-3)/(x^2-2x-3)] = (2x-3)/ (x^2+4x+3) [(x+2 )/{( x-3)(x+3)}] + [(x-3)/{(x-3)(x+1)}] = (2x-3)/ {(x+3)(x+1)} So the monomials appear in the denominator are x-3, x+3 and x+1. Make all the fractions have the same denominator: [(x+2 )/{( x-3)(x+3)}] + [(x-3)/{(x-3)(x+1)}] = (2x-3)/ {(x+3)(x+1)} [{(x+2 )(x+1)}/{( x-3)(x+3)(x+1)}] + [{(x-3)(x+3)}/{(x-3)(x+1)(x+3)}] = {(2x-3)(x-1)}/ {(x+3)(x+1)(x-1)} So the numerator we have (forget for a while the denominator) (x+2)(x+1) + (x-3)(x+3) = (2x-3)(x-1) (x^2 + 3x +2) + (x^2 - 9) = 2x^2 - 5x +3 8x = 10 x = 5/4

OpenStudy (anonymous):

wow you did the whole thing... I would've let vic23 do the basic algebric manipulations...

OpenStudy (anonymous):

THANK YOU VERY MUCH SKLEE ! :)

OpenStudy (anonymous):

Ya, I should leave some part for kathy to complete it ^_^

OpenStudy (anonymous):

sorry vic23 is not the guy who asked the question. My apologies. Let me correct myself. I would've let kathyworld09 do the basic algebric manipulations

OpenStudy (anonymous):

you're welcome kathy. Hope you learn something and understand it. I am sure you can do it next time :)

OpenStudy (anonymous):

so its not fully completed ?

OpenStudy (anonymous):

kathy, i have given the complete solution. :)

OpenStudy (anonymous):

You should learn and understand how it is done, so that you can do it again in other problem next time

OpenStudy (anonymous):

oh thank you , i just am having difficulty in a similar problem though